Exercise 8-1 Compute materials variances (L.O. 4)

During July, the cutting department competed 8,000 units of a product that had a standards materials cost of 2 square feet per unit at $2.40 per square foot. The actual materials purchased consisted of 16,400 square feet at $2.20 per square foot, for a total cost of $36,080. The actual material used this period was 16,160 square feet. Compute the materials price and usage variances. Indicated whether each is favorable or unfavorable. (Note: explain step by step , it should be clear and undrestandable).
Answer #1

Solution:

Standard quantity of material for actual production = 8000*2 = 16000 square feet

Standard price of material = $2.40 per foot

Actual quantity of material purchased = 16400 square feet

Actual quantity of material used = 16160 square feet

Actual price of material = $2.20 per square foot

Material price variance = (SP - AP) * AQ Purchased = ($2.40 - $2.20) * 16400 = $3,280 F

Materal usage variance = (SQ - AQ) * SP = (16000 - 16160) * $2.40 = $384 U